OpenStudy (anonymous):

**will give a medal** John began his job making $20 the first day. After that he was paid $6.00 per hour. Y = 6x + 20. Using the graph you drew for the equation, predict the pay for John if he works 7 hours after the first day.

OpenStudy (the_fizicx99):

y = 6x + 20 y = 6(7) + 20 42 + 20 y = 62 After 7 hours he will have earned $62

OpenStudy (anonymous):

thank you (:

OpenStudy (the_fizicx99):

Who gave you that function? Did you make it? f(x) = 6x + 20 ?

OpenStudy (anonymous):

no it was from a homework assignment lol

OpenStudy (the_fizicx99):

k